OpenStudy (anonymous):

what is the equation of line, L, through (-2,11) which is perpendicular to the line through the points (4,7) and (2,-6)? If the point (100,k) is on L, what is the value of k?

OpenStudy (gorv):

slope of perpendicular line =(7-(-6))/(4-2)=13/2 slope of line perpendicular to this =-2/13

OpenStudy (gorv):

y=mx+c y=-2x/13+c now put x=-2 and y=11 to find c